OpenStudy (anonymous):

Collect like terms: x^2+2xy+y^2=x^2-y^2

OpenStudy (anonymous):

x^2 terms cancel. 2xy + 2y^2 = 0 --> y(x+y) = 0. or yx + y^2 = 0.